OpenStudy (anonymous):

After 4 years, what is the total amount of an investment of $1500 in an account paying 2.5% interest, compounded quarterly? $2226.76 $1657.24 $1576.66 $1827.60

OpenStudy (anonymous):

wowee a whole different topic are you sure this is one class?

OpenStudy (anonymous):

paying 2.5% interest, compounded quarterly A= p(1+r/n) ^nt A = 1500(1 + .025/4)^16

OpenStudy (anonymous):

yes @satellite73

OpenStudy (anonymous):

what @jayshon_rogers314 wrote use this to compute http://www.wolframalpha.com/input/?i=1500%281+%2B+.025%2F4%29^16